OpenStudy (anonymous):

Simplify ((x/3)+(7/4))/((x/4)-(y/3)

OpenStudy (pfenn1):

\[\frac{\frac x3 + \frac 7 4}{\frac x4-\frac y3}\]Okay let's work with one term at a time. What is the LCD of \[\frac x3 +\frac 74\]

OpenStudy (anonymous):

idk

OpenStudy (eyust707):

what is the smallest number that 3 and 4 can both go in to?

OpenStudy (anonymous):

12

OpenStudy (eyust707):

correct!

OpenStudy (eyust707):

we call that number the least common denominator or LCD

OpenStudy (eyust707):

so for the first part the goal is to get a 12 in bottom of both fractions

OpenStudy (eyust707):

if we do that we can add them together

OpenStudy (eyust707):



OpenStudy (anonymous):

im sorry but im in a timed test i need to just have an answer and im vote you best answer and give you a medle please :3

OpenStudy (anonymous):

this is super important :(

OpenStudy (eyust707):

Im afraid that is in violation of the overall spirit of open study...

OpenStudy (eyust707):

I am more than willing to help you but answers I will not give.

OpenStudy (anonymous):

il do anything please! i need to finish this for my graduation and im running out of time :(

OpenStudy (eyust707):

you could have solved it by now...

OpenStudy (anonymous):

i cant thats why i asked open study

OpenStudy (eyust707):

if i multiply something by 1 does it change it?

OpenStudy (anonymous):

no

OpenStudy (eyust707):

correct! and would you agree that 4/4 = 1?

OpenStudy (anonymous):

yah

OpenStudy (eyust707):

great so check this out...

OpenStudy (eyust707):